Question 4

Perfect Furniture Berhad (PFB) specialises in living room leather furniture products. During the month of March, the company purchased RM210,000 of raw materials. Direct labour hours worked during the period were 8,000 hours. The workers were paid RM10 per hour. Direct expenses (high performance glue, etc) of RM22,000 were incurred. Manufacturing overhead for the month were RM72,000.

For the month of March, 72 units of leather furniture was sold for RM9,500 each. PFB also incurred selling costs including sales commission of RM155,000 and administration costs of RM32,000. Inventory information for the period is as follows:

1 March 2019

31 March 2019

Raw materials inventory

RM25,000

RM45,000

Work in process inventory

RM50,000

RM24,000

Finished goods inventory

RM10,400

RM26,000

Required:

  1. Determine the cost of raw materials used in production and prime costs.

                                                                                                                        (4 marks)

  1. Calculate the cost of goods manufactured for the month of March 2019.

                                                                                                                        (3 marks)

  1. Calculate the cost of goods sold for the month of March 2019.

                                                                                                                        (2 marks)

  1. Prepare a statement of profit or loss for the month of March 2019.

                                                                                                                        (3 marks)

                                                                                                           (Total: 12 Marks)

Question 5

Jokowi Bond Sdn Bhd (JBSB) produces a type of box which is sold for RM15 per unit. The normal annual production and sales for the boxes are 20,000 units.

The following data consist of costs incurred during the year ended 2018:

RM

Direct material

80,000

Direct Labour

50,000

Variable selling expenses

30,000

Administrative expenses (60% variable)

60,000

Fixed manufacturing overhead

20,000

The management accountant of the company is proposing the following alternatives to increase sales for the year 2019 and to reduce the idle capacity:

  1. Reducing the selling price to RM12 per unit which would lead to an estimated increase in the sales volume by 15%.
  2. An increase in sales would result in an increase of direct labour cost per unit by 10%.
  3. Fixed manufacturing overhead   is also expected to increase to RM25,000 due to an aggressive advertising campaign planned to boost sales.

Required:

  1. Calculate the total variable cost per unit and total fixed expenses for the year 2018. marks (4 marks)
  2. Calculate break even unit and sales for 2018.   marks (3 marks)          
  3. Calculate margin of safety in units and value for 2018                                                                                                                 (3 marks)
  4. Advise JBSB’s management on whether they should implement the proposal outlined above for the year 2019. (Show profit comparison).                                                                                                                                                  

(6 marks)

                                                                                     Question 4 Portor Furniture Band (PFBI speciale in ing room caller tumbare products. During the month of March the many purch
Answer #1
Please give positive ratings so I can keep answering. It would help me a lot. Please comment if you have any query. Thanks!
Perfect Furniture Berhad
Answer to Question 4
Raw materials transferred to production Amount RM
Beginning Raw material inventory      25,000.00
Add: Raw material Purchases 210,000.00
Total Raw material available for use 235,000.00
Less: Ending Raw material inventory      45,000.00
Cost of raw materials used in production 190,000.00
Calculation of Direct Labor costs Amount RM
Direct Labor hours worked        8,000.00
Rater per hour             10.00
Direct Labor costs     80,000.00
Prime costs Amount RM
Cost of raw materials used in production 190,000.00
Add: Direct Labor      80,000.00
Add: Direct Expenses      22,000.00
Prime costs 292,000.00
Cost of goods manufactured Amount RM
Cost of raw materials used in production 190,000.00
Direct Labor      80,000.00
Direct Expenses      22,000.00
Manufacturing Overhead      72,000.00
Total Manufacturing costs 364,000.00
Add: Beginning Work in process inventory      50,000.00
Total cost of work in process 414,000.00
Less: Ending Work in process inventory      24,000.00
Cost of goods manufactured 390,000.00
Cost of goods sold Amount RM
Beginning Finished goods inventory      10,400.00
Cost of goods manufactured 390,000.00
Cost of goods available for sale 400,400.00
Less: Finished goods inventory, ending      26,000.00
Cost of goods sold 374,400.00
Calculation of sales revenue Amount RM
Units sold             72.00
Price per unit        9,500.00
Sales revenue 684,000.00
Statement of profit or loss Amount RM
Sales Revenues
Sales Revenues 684,000.00
Net Sales 684,000.00
Less: Cost of goods sold 374,400.00
Gross Profit 309,600.00
Less: Operating Expenses
Sales commission 155,000.00
Administration costs      32,000.00
Total Operating Expenses 187,000.00
Net Income 122,600.00
Jokowi Bond Sdn Bhd
Answer to Question 5
Answer a
Calculation of total variable costs Amount RM Cost per unit Note
Number of units sold     20,000.00
Direct materials      80,000.00             4.00 This is $ 80,000 divided by 20,000 units.
Direct Labor      50,000.00             2.50 This is $ 50,000 divided by 20,000 units.
Variable selling expenses      30,000.00             1.50 This is $ 30,000 divided by 20,000 units.
Variable Administrative expenses      36,000.00             1.80 60% of Administrative expenses is variable. 60% of 60,000= $ 36,000. This is $ 36,000 divided by 20,000 units.
Total variable costs 196,000.00             9.80 A
Calculation of total fixed costs Amount RM Note
Fixed manufacturing overhead      20,000.00
Fixed Administrative expenses      24,000.00
Total fixed costs     44,000.00 B
Answer b Amount RM Note
Sell price per unit             15.00
Less: Variable cost per unit                9.80
Contribution per unit                5.20 C
Total fixed costs      44,000.00 See B
Breakeven point (units)        8,461.54 D=B/C
Sell price per unit             15.00 E
Breakeven point (sales) 126,923.08 F=D*E
Answer c Amount RM Note
Number of units sold      20,000.00 G
Less: Breakeven units        8,461.54 See D
Margin of safety in units     11,538.46 H=G-D
Amount RM
Number of units sold      20,000.00 See G
Sell price per unit             15.00 See E
Total sales 300,000.00 I=G*E
Less: Breakeven sales 126,923.08 See F
Margin of safety in value 173,076.92 J=I-F
Answer d
Revised situation Note
Units sold in 2018      20,000.00 See G
Increase in 2019 by 15%        3,000.00 K=G*15%
Units sold in 2019     23,000.00 L=G+K
Units sold in 2018      20,000.00 See G
Direct Labor      50,000.00 M
Direct Labor cost per unit                2.50 N=M/G
Increase in 2019 by 10%                0.25 O=N*10%
Direct Labor cost per unit in 2019                2.75 P=N+O
Comparative Income Statement 2018 2019 Note
Number of units sold     20,000.00 23,000.00 See L
Sell price             15.00           12.00 This is given in the question
Less: Variable costs per unit
Direct materials                4.00             4.00 See workings above
Direct Labor                2.50             2.75 See P
Variable selling expenses                1.50             1.50 See workings above
Variable Administrative expenses                1.80             1.80 See workings above
Total Variable costs per unit                9.80           10.05
Contribution per unit                5.20             1.95 Q
Total Contribution 104,000.00 44,850.00 R=P*Q
Less: Fixed costs
Fixed manufacturing overhead      20,000.00 25,000.00
Fixed Administrative expenses      24,000.00 24,000.00
Total fixed costs     44,000.00 49,000.00 S
Net Profit     60,000.00 (4,150.00) T=R-S
Decrease in Net Profit by 64,150.00 This is net profit of 2018 less net profit of 2019.
If the changes were done in 2019 then its net profit will decrease by $ 64,150. So the proposal should not be accepted.